# Learning to Ask: When LLM Agents Meet Unclear Instruction

### v3

@@ Abstract
Equipped with the capability to call functions, modern LLM agents can leverage external tools for addressing a range of tasks unattainable through language skills alone. However, the effective execution of these tools relies heavily not just on the advanced capabilities of LLM agents but also on precise user instructions, which often cannot be ensured in the real world. To evaluate the performance of LLM agents tool-use under imperfect instructions, we meticulously examine the real-world instructions queried from users, analyze the error patterns, and build a challenging tool-use benchmark called Noisy ToolBench (NoisyToolBench). We find that due to the next-token prediction training objective, LLM agents tend to arbitrarily generate the missed argument, which may lead to hallucinations and risks. To address this issue, we propose a novel framework, Ask-when-Needed (AwN), which prompts LLM agents to ask questions to users whenever they encounter obstacles due to unclear instructions. Moreover, to reduce the manual labor involved in user-LLM interaction and assess LLM agents’ performance in tool utilization from both accuracy and efficiency perspectives, we design an automated evaluation tool named ToolEvaluator. Our experiments demonstrate that the AwN significantly outperforms existing frameworks for tool learning in the NoisyToolBench. We will release all related code and datasets to support future research.
@@
Wenxuan Wang1  Juluan Shi1thanks:   Denote Equal Contribution.  Zixuan Ling1  Yuk-Kit Chan1  Chaozheng Wang1 Cheryl Lee1  Youliang Yuan2  Jen-tse Huang 1thanks:   Jen-tse Huang is the corresponding author.  Wenxiang Jiao2  Michael R. Lyu1 1The Chinese University of Hong Kong   2Tencent AI Lab 3The Chinese University of Hong Kong, Shenzhen 1{wxwang,jthuang,lyu}@cse.cuhk.edu.hk2joelwxjiao@tencent.com
@@ 1 Introduction
The incorporation of tool usage capabilities marks a pivotal step towards enhancing the intelligence of LLMs, pushing them closer to exhibiting human-like intelligence. The integration of tool usage allows LLMs to perform a broader array of complex and varied tasks, including managing emails, designing presentations, and browsing the web to gather real-time information. For example, LLMs can perform complex calculations using a calculator tool, access real-time weather updates through weather APIs, and execute programming code via interpreters Qin et al. (2023a); Schick et al. (2024); Mialon et al. (2023); Yang et al. (2023a). Toolformer Schick et al. (2024) is a pioneering work in empowering language models with self-learning capabilities for tool usage. Then, significant research efforts have been directed toward accessing a wider variety of tools or using multiple tools simultaneously to resolve a single query, such as GorillaPatil et al. (2023), RestGPT Song et al. (2023) and ToolLLM Qin et al. (2023b).
@@
To evaluate the performance of LLMs tool-use under unclear instruction, we design several innovative metrics from the accuracy and efficiency perspectives. For accuracy, we measure the LLMs’ proficiency in asking appropriate clarifying questions, their ability to execute the correct function calls, and their success in delivering final responses that meet the users’ needs. For efficiency, we calculate the average number of redundant asked questions and the average number of actions required to complete the instruction. An ideal LLM should achieve higher accuracy with fewer number of queries. Recognizing the labour-intensive nature of manually communicating with LLMs and verifying all execution results, we also innovatively design an automatic evaluation system, ToolEvaluator, to streamline the whole process. ToolEvaluator leverages the advanced problem-solving capabilities of GPT-4o to communicate with LLMs and automatically evaluate the performance of LLMs’ tool-using under unclear instruction. Our experiments on 6 LLMs and 2 tool-using frameworks demonstrate that the AwN significantly outperforms existing baseline methods for tool learning in the NoisyToolBench.
The key contributions of this research are summarized as follows:
@@ 2 Related Works
Tool Learning for LLMs. LLMs have recently made significant advancements, with ChatGPT being recognized as a major step towards achieving AGI Wu et al. (2023); Lund and Wang (2023); Jiao et al. (2023). These LLMs possess strong reasoning capabilities, enabling them to perform increasingly complex tasks Liu et al. (2023). However, to progress further towards AGI, it is crucial for LLMs to master the utilization of tools. Toolformer is the first innovative AI model designed to use several specialized tools, such as a web browser, a code interpreter, and a language translator, within a single framework Schick et al. (2023). The model’s ability to seamlessly switch between these tools and apply them contextually represents a significant advancement in AI capabilities. Recent studies like RestGPT Song et al. (2023) and ToolLLM Qin et al. (2023b), have connected LLMs with real-life Application Programming Interfaces (APIs), allowing them to sequentially employ multiple external tools to solve user queries. The tool-augmented approach empowers LLMs to use various kinds of tools to do more sophisticated tasks, showcasing an enhanced level of capability compared to pure LLMs. Besides, API-Bank Li et al. (2023), ToolAlpaca Tang et al. (2023), ToolBench Qin et al. (2023b), ToolQA Zhuang et al. (2023) and RestBench Song et al. (2023) are exemplary benchmarks to systematically evaluate the performance of tool-augmented LLMs performance in response to user’s queries. However, current models often ignore the situations in which users might not give exact instructions, which can result in the tools not working properly. Thus, our study aims to tackle this specific challenge by developing a new benchmark specifically for ambiguous instructions.
@@
Learning to Ask. Since user queries may not always be clear, and the execution of LLMs may encounter uncertainties and ambiguities, learning to ask questions has emerged as a challenging yet crucial research area Rao and Daumé III (2018); Kuhn et al. (2022); Andukuri et al. (2024). For example, some researchers introduce a learning framework that empowers an embodied visual navigation agent to proactively seek assistanceZhang et al. (2023). Recently, similar ideas have been adopted in the software engineering, leveraging a communicator to enhance the reliability and quality of generated code Wu (2023). Our work focuses on the tool-learning scenario, which is more sensitive to the user’s unclear query. A concurrent study Qian et al. (2024) also focuses on the reliability of tool-learning systems under unclear instruction. However, they did not systematically examine real-world user behavior, leading to the limited and biased nature of their dataset that doesn’t accurately capture user errors. Our research addresses this shortfall by starting with a user analysis. Additionally, Qian’s methodology depends significantly on human manual interaction and assessment of LLM performances, which is time-consuming and hard to reproduce. In contrast, we introduce an automated evaluation method that can proxy humans to communicate with and automatically evaluate the performance of LLMs.
@@ 5.1 Evaluation Metrics
Model Framework IMKI IMR IwE IBTC A1(%) A2(%) A3(%) A1(%) A2(%) A3(%) A1(%) A2(%) A3(%) A1(%) gpt-3.5 CoT 0.74 0.36 0.22 0.20 0.24 0.12 0.5 0.24 0.16 0.38 + AwN 0.74 0.44 0.24 0.86 0.46 0.20 0.74 0.48 0.28 0.48 DFSDT 0.64 0.16 0.12 0.60 0.18 0.16 0.48 0.14 0.14 0.46 + AwN 0.88 0.52 0.46 0.88 0.56 0.48 0.72 0.42 0.36 0.64 gpt-4 CoT 0.74 0.48 0.32 0.72 0.52 0.36 0.52 0.26 0.24 0.34 + AwN 0.94 0.62 0.50 0.76 0.44 0.38 0.48 0.34 0.34 0.94 DFSDT 0.82 0.16 0.16 0.70 0.28 0.26 0.54 0.12 0.10 0.54 + AwN 0.80 0.56 0.48 0.80 0.50 0.44 0.52 0.38 0.36 0.94 gpt-4o CoT 0.52 0.48 0.34 0.18 0.28 0.16 0.12 0.12 0.10 0.10 + AwN 0.90 0.58 0.36 0.80 0.46 0.30 0.60 0.44 0.32 0.92 DFSDT 0.58 0.20 0.18 0.26 0.18 0.16 0.18 0.06 0.04 0.08 + AwN 0.88 0.60 0.46 0.90 0.52 0.36 0.64 0.46 0.38 0.94 deepseek-v3 CoT 0.44 0.40 0.20 0.24 0.28 0.24 0.10 0.14 0.14 0.30 + AwN 0.70 0.52 0.36 0.70 0.54 0.46 0.40 0.30 0.26 0.98 DFSDT 0.42 0.30 0.26 0.60 0.20 0.18 0.22 0.12 0.12 0.48 + AwN 0.72 0.52 0.42 0.82 0.52 0.48 0.54 0.38 0.36 0.98 gemini-1.5 CoT 0.22 0.18 0.10 0.22 0.10 0.02 0.08 0.12 0.06 0.52 + AwN 0.86 0.40 0.18 0.74 0.24 0.08 0.58 0.28 0.22 0.68 DFSDT 0.62 0.02 0.02 0.6 0.08 0.04 0.36 0.06 0.02 0.48 + AwN 0.82 0.40 0.12 0.76 0.28 0.04 0.66 0.36 0.26 0.70 claude-3.5 CoT 0.24 0.26 0.20 0.12 0.28 0.24 0.08 0.26 0.24 0.30 + AwN 0.54 0.5 0.5 0.32 0.30 0.24 0.34 0.34 0.26 0.88 DFSDT 0.26 0.18 0.14 0.12 0.18 0.18 0.12 0.20 0.18 0.62 + AwN 0.52 0.44 0.42 0.32 0.30 0.18 0.36 0.36 0.30 0.86
@@ 5.2 Auto-Evaluation Pipeline
Auto-Assessment. ToolEvaluator can also automatically assess how well LLMs perform under ambiguous instructions according to the five metrics introduced above. A1 measures whether LLMs can ask the right question. ToolEvaluator calculates the semantic similarity between the LLMs-asked question and the expected question to asses A1. A2 measures whether LLMs can conduct correct API calls. Following the previous works Yang et al. (2023b); Chiang and yi Lee (2023); Wang et al. (2023); Yuan et al. (2023), ToolEvaluator adopts GPT-4o as a judge to identify whether the generated API calls are the same as the expected API calls. A3 measures whether LLMs can correctly generate the final answer. ToolEvaluator adopts GPT-4o as a judge to identify whether the final answer aligns with the user’s intent. For measuring the efficiency, ToolEvaluator counts the number of generated irrelevant questions as Re and counts the total number of actions during the process as Steps.
@@ 5.5 Main Result
AwN can reduce the average cost of LLM’s tool-using. The average number of steps measures the cost of LLMs’ tool-using. As is shown in Table 3, adopting AwN can reduce the number of actions, especially for gpt-4-based DFSDT, deepseek-v3-based-DFSDT and gemini-1.5-based-DFSDT. Although AwN can lead to a higher cost for a few LLM agents, such as claude-3.5, considering the significant performance improvements achieved, the moderate increase in cost is justifiable and worthwhile.
@@ 6 Conclusion
This study explores how unclear user instructions hinder modern LLMs’ tool usage. To investigating the common error patterns in real-world instructions, we propose: (1) Noisy ToolBench (NoisyToolBench), a novel benchmark for evaluating LLM performance under ambiguous instructions; (2) Ask-when-Needed (AwN), an innovative approach enabling LLMs to request clarification when uncertain; and (3) an automated evaluator (ToolEvaluator) to assess accuracy and efficiency. Experimental results show that the AwN algorithm markedly surpasses existing methods in the NoisyToolBench dataset and significantly improves the performance of LLMs’ tool-using under unclear user instructions.

### v4

@@ Abstract
Equipped with the capability to call functions, modern LLM agents can leverage external tools for addressing a range of tasks unattainable through language skills alone. However, the effective execution of these tools relies heavily not just on the advanced capabilities of LLM agents but also on precise user instructions, which often cannot be ensured in the real world. To evaluate the performance of LLM agents tool-use under imperfect instructions, we meticulously examine the real-world instructions queried from users, analyze the error patterns, and build a challenging tool-use benchmark called Noisy ToolBench (NoisyToolBench). We find that due to the next-token prediction training objective, LLM agents tend to arbitrarily generate the missed argument, which may lead to hallucinations and risks. To address this issue, we propose a novel framework, Ask-when-Needed (AwN), which prompts LLM agents to ask questions to users whenever they encounter obstacles due to unclear instructions. Moreover, to reduce the manual labor involved in user-LLM interaction and assess LLM agents’ performance in tool utilization from both accuracy and efficiency perspectives, we design an automated evaluation tool named ToolEvaluator. Our experiments demonstrate that the AwN significantly outperforms existing frameworks for tool learning in the NoisyToolBench. We release all related code and datasets to support future research [^1] https://github.com/Mysterchan/learning_to_ask.
@@
Wenxuan Wang1   Juluan Shi2thanks:   Denote Equal Contribution.   Zixuan Ling2   Yuk-Kit Chan2   Chaozheng Wang2 Cheryl Lee2   Youliang Yuan3   Jen-tse Huang 4thanks:   Jen-tse and Wenxiang are corresponding authors.   Wenxiang Jiao5†   Michael R. Lyu2 1Renmin University of China   2The Chinese University of Hong Kong 3The Chinese University of Hong Kong, Shenzhen 4Johns Hopkins University  5Xiaohongshu Inc. 1wangwenxuan@ruc.edu.cn   4jhuan236@jh.edu   5wenxiangjiaonju@gmail.com
@@ 1 Introduction
The incorporation of tool usage capabilities marks a pivotal step towards enhancing the intelligence of LLMs, pushing them closer to exhibiting human-like intelligence. The integration of tool usage allows LLMs to perform a broader array of complex and varied tasks. For example, LLMs can perform complex calculations using a calculator tool, access real-time weather updates through weather APIs, and execute programming code via interpreters Qin et al. (2023a); Schick et al. (2024); Mialon et al. (2023); Yang et al. (2023a). Toolformer Schick et al. (2024) is a pioneering work in empowering language models with self-learning capabilities for tool usage. Then, significant research efforts have been directed toward accessing a wider variety of tools or using multiple tools simultaneously to resolve a single query, such as GorillaPatil et al. (2023), RestGPT Song et al. (2023) and ToolLLM Qin et al. (2023b).
@@
To evaluate the performance of LLMs tool-use under unclear instruction, we design several innovative metrics from the accuracy and efficiency perspectives. For accuracy, we measure the LLMs’ proficiency in asking appropriate clarifying questions, their ability to execute the correct function calls, and their success in delivering final responses that meet the users’ needs. For efficiency, we calculate the average number of redundant asked questions and the average number of actions required to complete the instruction. An ideal LLM should achieve higher accuracy with fewer number of queries. Recognizing the labour-intensive nature of manually communicating with LLMs and verifying all execution results, we also innovatively design an automatic evaluation system, ToolEvaluator, to streamline the whole process. ToolEvaluator leverages the advanced problem-solving capabilities of GPT-4o to communicate with LLMs and automatically evaluate the performance of LLMs’ tool-using under unclear instructions. Our experiments on 8 LLMs and 2 tool-using frameworks demonstrate that the AwN significantly outperforms existing baseline methods.
The contributions are summarized as follows:
@@ 2 Related Works
Tool Learning for LLMs. LLMs have recently made significant advancements, with ChatGPT being recognized as a major step towards achieving AGI Wu et al. (2023); Lund and Wang (2023); Jiao et al. (2023). However, to progress further towards AGI, it is crucial for LLMs to master the utilization of tools. Toolformer is the first innovative AI model designed to use several specialized tools, such as a web browser, a code interpreter, and a language translator, within a single framework Schick et al. (2023). The model’s ability to seamlessly switch between these tools and apply them contextually represents a significant advancement in AI capabilities. Recent studies like RestGPT Song et al. (2023) and ToolLLM Qin et al. (2023b), have connected LLMs with real-life Application Programming Interfaces (APIs), allowing them to sequentially employ multiple external tools to solve user queries. The tool-augmented approach empowers LLMs to use various kinds of tools to do more sophisticated tasks, showcasing an enhanced level of capability compared to pure LLMs. Besides, API-Bank Li et al. (2023), ToolAlpaca Tang et al. (2023), ToolBench Qin et al. (2023b), ToolQA Zhuang et al. (2023) and RestBench Song et al. (2023) are exemplary benchmarks to systematically evaluate the performance of tool-augmented LLMs performance in response to user’s queries. However, current models often ignore the situations in which users might not give exact instructions, which can result in the tools not working properly. Thus, our study aims to tackle this specific challenge by developing a new benchmark specifically for ambiguous instructions.
@@
Learning to Ask. Since user queries may not always be clear, and the execution of LLMs may encounter uncertainties and ambiguities, learning to ask questions has emerged as a challenging yet crucial research area Rao and Daumé III (2018); Kuhn et al. (2022); Andukuri et al. (2024). For example, some researchers introduce a learning framework that empowers an embodied visual navigation agent to proactively seek assistanceZhang et al. (2023). Recently, similar ideas have been adopted in the software engineering, leveraging a communicator to enhance the reliability and quality of generated code Wu (2023). Our work focuses on the tool-learning scenario, which is more sensitive to the user’s unclear query. A concurrent study Qian et al. (2024) also focuses on the reliability of tool-learning systems under unclear instruction. However, they did not systematically examine real-world user behavior, leading to the limited and biased nature of their dataset that doesn’t accurately capture user errors. Additionally, Qian’s methodology depends significantly on human manual interaction and assessment of LLM performances, which is time-consuming and hard to reproduce.
@@ 5.1 Evaluation Metrics
Model Framework IMKI IMR IwE IBTC A1(%) A2(%) A3(%) A1(%) A2(%) A3(%) A1(%) A2(%) A3(%) A1(%) gpt-3.5 CoT 0.74 0.36 0.22 0.20 0.24 0.12 0.5 0.24 0.16 0.38 + AwN 0.74 0.44 0.24 0.86 0.46 0.20 0.74 0.48 0.28 0.48 DFSDT 0.64 0.16 0.12 0.60 0.18 0.16 0.48 0.14 0.14 0.46 + AwN 0.88 0.52 0.46 0.88 0.56 0.48 0.72 0.42 0.36 0.64 gpt-4 CoT 0.74 0.48 0.32 0.72 0.52 0.36 0.52 0.26 0.24 0.34 + AwN 0.94 0.62 0.50 0.76 0.44 0.38 0.48 0.34 0.34 0.94 DFSDT 0.82 0.16 0.16 0.70 0.28 0.26 0.54 0.12 0.10 0.54 + AwN 0.80 0.56 0.48 0.80 0.50 0.44 0.52 0.38 0.36 0.94 gpt-4o CoT 0.52 0.48 0.34 0.18 0.28 0.16 0.12 0.12 0.10 0.10 + AwN 0.90 0.58 0.36 0.80 0.46 0.30 0.60 0.44 0.32 0.92 DFSDT 0.58 0.20 0.18 0.26 0.18 0.16 0.18 0.06 0.04 0.08 + AwN 0.88 0.60 0.46 0.90 0.52 0.36 0.64 0.46 0.38 0.94 deepseek-v3 CoT 0.44 0.40 0.20 0.24 0.28 0.24 0.10 0.14 0.14 0.30 + AwN 0.70 0.52 0.36 0.70 0.54 0.46 0.40 0.30 0.26 0.98 DFSDT 0.42 0.30 0.26 0.60 0.20 0.18 0.22 0.12 0.12 0.48 + AwN 0.72 0.52 0.42 0.82 0.52 0.48 0.54 0.38 0.36 0.98 gemini-1.5 CoT 0.22 0.18 0.10 0.22 0.10 0.02 0.08 0.12 0.06 0.52 + AwN 0.86 0.40 0.18 0.74 0.24 0.08 0.58 0.28 0.22 0.68 DFSDT 0.62 0.02 0.02 0.6 0.08 0.04 0.36 0.06 0.02 0.48 + AwN 0.82 0.40 0.12 0.76 0.28 0.04 0.66 0.36 0.26 0.70 claude-3.5 CoT 0.24 0.26 0.20 0.12 0.28 0.24 0.08 0.26 0.24 0.30 + AwN 0.54 0.5 0.5 0.32 0.30 0.24 0.34 0.34 0.26 0.88 DFSDT 0.26 0.18 0.14 0.12 0.18 0.18 0.12 0.20 0.18 0.62 + AwN 0.52 0.44 0.42 0.32 0.30 0.18 0.36 0.36 0.30 0.86 O3-mini CoT 0.00 0.16 0.16 0.00 0.16 0.14 0.00 0.04 0.04 0.00 + AwN 0.78 0.54 0.52 0.76 0.54 0.48 0.58 0.42 0.32 0.40 DFSDT 0.00 0.16 0.16 0.00 0.16 0.10 0.00 0.06 0.02 0.00 + AwN 0.80 0.54 0.52 0.76 0.58 0.54 0.58 0.38 0.36 0.48 DeepSeek-R1 CoT 0.10 0.22 0.16 0.02 0.28 0.12 0.00 0.08 0.06 0.00 + AwN 0.80 0.52 0.34 0.60 0.54 0.22 0.50 0.34 0.22 0.84 DFSDT 0.02 0.20 0.18 0.04 0.26 0.14 0.00 0.08 0.04 0.00 + AwN 0.86 0.62 0.42 0.62 0.42 0.26 0.54 0.38 0.22 0.76
@@ 5.2 Auto-Evaluation Pipeline
Auto-Assessment. ToolEvaluator can also automatically assess how well LLMs perform under ambiguous instructions according to the five metrics introduced above. A1 measures whether LLMs can ask the right question. ToolEvaluator calculates the semantic similarity between the LLMs-asked question and the expected question to asses A1. A2 measures whether LLMs can conduct correct API calls. Following the previous works Yang et al. (2023b); Chiang and Lee (2023); Wang et al. (2023); Yuan et al. (2023), ToolEvaluator adopts GPT-4o as a judge to identify whether the generated API calls are the same as the expected API calls. A3 measures whether LLMs can correctly generate the final answer. ToolEvaluator adopts GPT-4o as a judge to identify whether the final answer aligns with the user’s intent. For measuring the efficiency, ToolEvaluator counts the number of generated irrelevant questions as Re and counts the total number of actions during the process as Steps.
@@ 5.5 Main Result
AwN can also improve the performance of reasoning models. Experimental results on two recently released strong reasoning models, OpenAI O3-mini and DeepSeek R1, showing that 1) vanilla models still cannot achieve good performance under unclear user instruction; 2) although vanilla reasoning models do not ask questions, it can sometimes produce the correct function call (A2 > A1), due to their powerful reasoning abilities; 3) AwN can significantly improve the performance by prompting LLMs to ask good questions.
@@
AwN can reduce the average cost of LLM’s tool-using. The average number of steps measures the cost of LLMs’ tool-using. As is shown in Table 3, adopting AwN can reduce the number of actions, especially for gpt-4-based DFSDT, deepseek-v3-based-DFSDT and gemini-1.5-based-DFSDT. Although AwN can lead to a higher cost for a few LLM agents, such as claude-3.5, considering the significant performance improvements achieved, the moderate increase in cost is justifiable.
6 Discussion
Our study provides several key insights into addressing unclear user instructions in LLM-based tool use.
First, the taxonomy of four ambiguity types—Instructions Missing Key Information (IMKI), Instructions with Multiple References (IMR), Instructions with Errors (IwE), and Instructions Beyond Tool Capabilities (IBTC)—derived from analyzing 1,000 real-world user queries, offers a robust framework for understanding prevalent instruction challenges. Identified through manual annotation by 10 qualified annotators with a two-round cross-verification process, these categories ensure a realistic and controlled evaluation of LLM performance under imperfect instructions. While real-world ambiguities can involve more complex, context-dependent, or cumulative issues in multi-step tasks, our taxonomy, as shown in Table 1 (e.g., 56.0% IMKI), captures dominant issues and provides a foundational step for future research into intricate scenarios.
Second, the AwN framework, though a simple prompt-engineering approach, demonstrates significant effectiveness in improving LLM tool-use under unclear instructions. By prompting LLMs to proactively seek clarification, AwN markedly enhances accuracy across various models (e.g., A1 scores for gpt-4o-based CoT improved from 0.52 to 0.90, Table 2). Its compatibility with diverse LLMs and frameworks like CoT and DFSDT, combined with a 4.2/5 user satisfaction score and 82.5% resolution rate for unclear queries (Appendix A.3), highlights its practical utility. The simplicity of AwN is a strength, offering a scalable baseline for practitioners. Future work could explore fine-tuning LLMs on NoisyToolBench or integrating AwN with advanced agent frameworks to further enhance performance under ambiguous instructions.
Finally, NoisyToolBench’s focus on unclear instructions is a deliberate design choice to isolate and evaluate LLM performance in challenging scenarios. In real-world settings, LLMs must discern instruction clarity, but our controlled benchmark allows precise assessment of their ability to handle ambiguities. Experiments on clear instructions (Appendix A.1, Table 4 and Table 5) confirm that AwN maintains performance in such cases, suggesting its robustness across mixed scenarios. Future work could integrate clear and unclear instructions to test LLMs’ discernment capabilities, building on AwN’s foundation to address more dynamic real-world interactions.
@@ 7 Conclusion
This paper explores how unclear user instructions hinder LLM agents’ tool usage by proposing: (1) Noisy ToolBench (NoisyToolBench), a novel benchmark for evaluating LLM performance under ambiguous instructions; (2) Ask-when-Needed (AwN), an innovative approach enabling LLMs to request clarification when uncertain; and (3) an automated evaluator (ToolEvaluator) to assess accuracy and efficiency. Experimental results show that the AwN algorithm significantly improves the performance of LLMs’ tool-using under unclear user instructions.
@@
Ethics Statement
This paper reports a user experience study conducted to evaluate the effectiveness of AwN. All participants were informed in advance that their feedback would be used for research purposes and might appear in this publication. No personal or sensitive data were collected, and all responses were anonymized to protect privacy. Participation was entirely voluntary, and volunteers were selected based on their proficiency with LLMs to ensure informed and meaningful contributions. Our methodology follows responsible research practices, emphasizing transparency, fairness, and respect for participants.
Appendix A Appendix
A.1 AwN Does Not Affect the Performance on Clear Instructions
Since NoisyToolBench focuses on evaluating the performance of LLM tool-use agents under unclear user instructions, the data only contains unclear instructions, which can complement existing tool-use evaluation datasets to assess the performance under imperfect user instructions. However, evaluating AwN on clear instructions is also important. So we select 200 samples from ToolBench, the user queries of which are clear, and we evaluate the performance of our Awn on this set. As is shown in Table 4, AwN can significantly improve the performance on unclear instructions and does not affect the performance under clear instructions.
To evaluate the generalizability of AwN, we tested it on a 200-query subset of ShortcutsBench Shen et al. (2024), a benchmark featuring complex real-world API-based tasks with clear instructions. We measured API selection accuracy (correctly choosing the relevant API) and API calling accuracy (executing the API with correct arguments). Table 5 shows that AwN consistently improves API selection accuracy (e.g., +6.5% for GPT-4o) and maintains or enhances API calling accuracy (e.g., +4.0% for Claude) across all models. These results demonstrate that AwN not only excels with ambiguous instructions but also generalizes effectively to clear, real-world tool-use scenarios.
Model Clean Noisy_IMKI Noisy_IMR Noisy_IwE
GPT-4o 0.54 0.34 0.16 0.10
+ AwN 0.51 0.36 0.30 0.32
DeepSeek V3 0.64 0.20 0.24 0.14
+ AwN 0.64 0.36 0.46 0.26
Claude 3.5 0.60 0.20 0.24 0.24
+ AwN 0.64 0.50 0.30 0.26
Gemini 1.5 0.36 0.10 0.02 0.06
+ AwN 0.40 0.18 0.08 0.22
Table 4: Assessing the performance of AwN on both clear and unclear instructions
Method API Selection Acc. API Call Acc.
GPT-4o 86.0 45.0
+ AwN 92.5 42.5
DeepSeek V3 89.0 53.5
+ AwN 91.5 53.5
Claude 3.5 89.0 46.0
+ AwN 88.5 50.0
Gemini 1.5 83.5 46.0
+ AwN 87.0 46.0
Table 5: Performance of AwN on ShortcutsBench with clear instructions (200 queries).
A.2 Details of the Calculation of Each Metric
ToolEvaluator evaluated the model performance on A1, Re and Steps by a deterministic, rule-based algorithm that analyzes the responses. A2 and A3 are evaluated by LLM-as-a-judge.
The following are the details of the calculation of each metric.
A1: A1 is computed by analyzing whether the LLMs generate valid clarifying questions under ambiguous instructions in the NoisyToolbench dataset. A binary flag is given for each instruction; it is set to 1 if there is a clarifying question identified by similarity check. Otherwise, it remains 0.
Re: Re is computed by counting the number of redundant questions asked by LLMs. Similar to A1 situations, each clarifying question raised by LLM will be classified into relevant questions for ambiguity resolution or irrelevant questions. Re here is equal to the number of mis-clarifying questions raised by LLMs under each instruction.
Steps: Steps are computed by counting the number of total actions (generate response, ask question and call API) performed by LLM. The algorithm will go through the whole conversation history for each instruction.
A2 and A3: ToolEvaluator adopts LLMs as evaluators. We adopt GPT-4o as the evaluator to score A2 and A3 for objective 1 (whether using the correct API) and objective 2 (whether returning the correct answer). The instruction is:
As an evaluator of tool-augmented language model systems, your responsibility is to assess the models’ effectiveness in using APIs to gather the necessary information to fulfill user requests. This involves reviewing the actual API calls made by the models against a given set of required API calls, including their parameters. Your evaluation focuses on two main objectives:
Objective 1: Verify the accuracy of the actual tool calls made by the models against the expected tool calls, including their arguments. A successful outcome means that the models executed all required API calls with expected arguments. Then the score of the Objective 1 should be 1. Otherwise, it is a failure for Objective 1 and the score should be 0. Here are some examples …
Objective 2: Assess whether the model’s final response correctly achieve the user’s instruction and check if the final answer was indeed based on the data retrieved from these API calls, as opposed to being generated independently of these tool calls. Success in this objective means the model effectively used the API calls to achieve user’s instructions. Conversely, failure suggests the response was not derived from the API data or the user’s instruction is not achieved. It’s important to note that any thoughts of the LLMs not based on the API response, especially regarding makeup information, are not considered valid answers. Please note that if the model does not provide the final answer, we consider it as a failure case. Here are some examples …
A.3 User Experiments
To show the effectiveness of our method, we follow Wang et al. (2024) to conduct a user experience study with real-world users. Specifically, we recruit 5 volunteers in Section 3.1 back, who have a Bachelor’s degree, are proficient in English, and have experience using LLMs. We provided them with 10 APIs, along with the user instructions they designed to call the APIs, as well as the feedback AwN generated. Then we ask the following questions: From 1 (very disagree) to 5 (strongly agree), how much do you agree that the system’s clarifying questions are relevant and helpful to solving the request? We got an average score of 4.2, showing the effectiveness of AwN from the real-world user perspectives.